classify the conic section defined by the equation 4x^2+9x^2-8x-18y-23=0
mulitple choice answers
hyparbola
circle
parabola
ellispe Answer by lwsshak3(11628) 
 classify the conic section defined by the equation 4x^2+9x^2-8x-18y-23=0
**
My guess you meant 9x^2 to be 9y^2, so I will assume this change in my solution
4x^2+9y^2-8x-18y-23=0
Complete the square
4(x^2-4x+4)+9(y^2-2y+1)=23+16+9
4(x-2)+9(y-1)^2=48
(x-2)/12+9(y-1)^2/(16/3)=1
This is an equation of an ellipse with horizontal major axis